OpenStudy (allie2020):

An angle that shares the same sine value of an angle that measures 5 pi/4 radians is located where? Quadrant I Quadrant II Quadrant IV along an axis

OpenStudy (allie2020):

@jtug6 @Mehek14

jhonyy9 (jhonyy9):

so do you know that 5pi/4 = 5*180/4 = 900/4 = 225 degrees 225-180 = 45 degrees and will be located in quadrant III